Current is applied to a molten mixture of pbF_2, F What is produced at cathode and anode?
Answer #1

Eo (Pb2+/Pb) = -0.13 V
Eo(Fe2+/Fe) = -0.44 V
Eo (Al3+/Al)= -1.66 V
Eo(F2/F-)= 2.87 V
Eo(Cl2/Cl-)= 1.36 V
Eo(Br2/Br-)=1.09 V

the one with greateest Eo value will be reduced at cathode
The Eo with the smallest EO value will be oxidised at anode

At cathode:
F2 is produced

at anode: Al is produced